The ages of commercial aircraft are normally distributed with a mean of 13 years and a standard deviation of 7.7672 years . what percentage of individual aircraft have ages between nine years and 16 years ? assume that a random sample of 64 aircraft is selected and the mean age of the sample is computed what percentage of sample means had ages between nine years and 16 years?
Answer by Boreal(15235)

this is a z test
for first part, z=(x-mean)/sd, and that is between -4/7.7672 or -0.51 and z=3/7.7672 or +0.39. That probability is 0.3467

z=(x bar- mean)/sd/sqrt(n)
lower z is > (9-13)*8/7.7672, moving sqrt (64)=8 to numerator.
That is z >-4.12
upper z is 3*8/7.7672
that is z>3.09
that probability with those z constraints is 0.9990.